OpenStudy (anonymous):

Getting another Calculator Error Cosine Law Trig Here is the problem A = 103 degrees Sides are b =31 c=52 Equation \( a^2 = 31^2+52^2-2\times 31 \times 52 \) = a = 66.2589 \( \large \cos B = \frac{66.25^2 +52^2-31^2}{2\times 66.25\times52} \) Am I right so far?? Do you get an error?

OpenStudy (abdullahm):

Where is cos(A) in the equation?

OpenStudy (abdullahm):

\( \sf a^2 = 31^2+52^2-2\times 31 \times 52 cos(103)\) = 66.2589

OpenStudy (abdullahm):

Oh nevermind, you forgot to put cos(103) but you arrived at the correct a value.

OpenStudy (anonymous):

Did I forget to square that??

OpenStudy (abdullahm):

\( \sf a^2 = 31^2+52^2-2\times 31 \times 52 cos(103)\) then a=66.2589 not a^2, my bad.

OpenStudy (anonymous):

Ok cool. Do you get an error on your cal??

OpenStudy (anonymous):

\( \huge B =\cos^{-1}( \frac{66.25^2 +52^2-31^2}{2\times 66.25\times52} )\) I get an error

OpenStudy (anonymous):

Crap it worked this time........ I got 27.12

OpenStudy (abdullahm):

So now we have: A= 103 degrees a = 66.26 b=31 c=52 Using sin laws \(\sf\Large \frac{sin103}{66.26} =\frac{sinB}{31}=\frac{sinC}{52}\) So we want to solve for \(\sf B\). \(\sf\Large \frac{sin103}{66.26} =\frac{sinB}{31}\) Cross multiply \(\sf\Large sin(103) \times 31 = Sin B\times 66.26\)

OpenStudy (abdullahm):

Yes, 27.12 is the correct value for B :)

OpenStudy (anonymous):

That is strange..... I did not do anything . I just hit my back button to bring it up again and then hit enter and it gave me the answer. LOL

OpenStudy (anonymous):

Wierd.... Stuff is going to drive me batty

OpenStudy (abdullahm):

Calculators are evil lol

OpenStudy (anonymous):

Thanks for the help. :-)

OpenStudy (abdullahm):

Anytime :)

OpenStudy (abdullahm):

Once you get the hang of the questions, it will become more easier xD